Circular No. 5722

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

U GEMINORUM J. Mattei, AAVSO, reports that this cataclysmic variable is undergo-

ing an outburst, as indicated by the following visual magnitude estimates: Mar. 12.03 UT, 14.6 (J. Bortle, Stormville, NY); 16.78, 14.4 (R. Fidrich, Ibafa, Hungary); 18.84, 13.3 (G. Poyner, Birmingham, England); 19.07, 13.4 (Bortle); 19.20, 12.9 (B. Stewart, Rochelle Park, NJ); 19.32, 11.2 (R. Royer, Lakewood, CA); 19.54, 11.1 (Royer); 19.71, 9.0 (O. Gabzo, Kibbutz Sheffagim, Israel); 19.82, 9.5 (K. Medway, Southampton, England). The last outburst of U Gem was in Aug. 1992, when it reached mag 9.3 on Aug. 30 and was brighter than mag 11 for 12 days. U Gem has been monitored extremely closely since, and no other outburst activity has been recorded.

SUPERNOVA 1993G IN NGC 3690 D. D. Balam, L. Bogan, D. Zurek and R. M. Robb report a measure-

ment of V = 17.2 ± 0.2 on Mar. 16.42 UT, using the 0.5-m reflector (+ CCD) of the University of Victoria’s Climenhaga Observatory.

S. Pesci, Milan, Italy, reports that the supernova was not present to limiting mag 15.6 on Feb. 11.8 and 19.8 UT.

Corrigendum. On IAUC 5720, the photospheric expansion velocity should have been given as ∼ 6000 km/s, not ∼ 600 km/s.

29999-231-9304 Subscribers to these Circulars are reminded that the last four digits of

the 12-figure number on the address labels indicate the year and month of the account’s expiry date. The above example means that no Circulars will be sent after the end of Apr. 1993. This information is particularly necessary for ‘special’ subscribers, who do not receive invoices or other reminders to renew; to guarantee no interruption in service, renewal payments ($6.00 per month for the printed edition alone, $13.50 per month to include e-mail delivery) should be sent to the postal address given above early enough that it is recorded when the accounts are updated at the end of the previous month. Please note your account number (the first five digits of the 12-figure number) when making payments (which should normally be for 6, 12 or 18 months).

1993 March 21 Brian G. Marsden